OpenStudy (rational):

find the period of \(\large 1/83 \) in its decimal expansion

myininaya (myininaya):

what is period mean here?

myininaya (myininaya):

is that just any repeating part that might occur in the expansion?

OpenStudy (rational):

period is the length of the repeating part... example : the period of 1/37 is 3 because 1/37 = 0.027027027...

myininaya (myininaya):

the length of the repeating part

OpenStudy (mathmath333):

82

myininaya (myininaya):

This one is going to be pretty tough without any tricks because it it is suppose to be 41. :p

myininaya (myininaya):

how did you get 82

OpenStudy (mathmath333):

\((10^{41}-1) \pmod {83}=0\)

OpenStudy (rational):

yes both 41 and 82 are periods.. but 41 is the smallest one actually i think we can say any multiple of 41 is a period

OpenStudy (mathmath333):

also \((10^{82}-1) \pmod {83}=0\)

OpenStudy (mathmath333):

fermat little theorm

OpenStudy (rational):

interesting, whats that and does that work

OpenStudy (mathmath333):

and how did u get 41 @myininaya

OpenStudy (rational):

Ohh I see.. how is that related to period ?

OpenStudy (mathmath333):

found here http://hr.userweb.mwn.de/numb/period.html the length of period of the denominator \(q\) should divide \(10^{k}-1\)

myininaya (myininaya):

lol I was trying to find a trick how to get 41

OpenStudy (mathmath333):

but sometimes \(\Large m^{\phi (n)} =m^{\phi(n)/2} \equiv 1 \pmod q\)

OpenStudy (mathmath333):

\(\Large m^{\phi (n)} =m^{\phi(n)/2} \equiv 1 \pmod n\)

myininaya (myininaya):

\[(10^{x}-1) \pmod {n}=0\] so finding x here will give us the period for 1/n ---would this work for all n and x

OpenStudy (mathmath333):

no u have to check the earlier powers to specially its \(\phi(n)/2\)

OpenStudy (rational):

n = 0.abcdef abcdef abcdef ... 10^6n=abcdef.abcdef abcdef abcdef ... -------------------------------------- 10^6n-n = abcdef n(10^6-1) = abcdef yeah it seem to work in general xD

myininaya (myininaya):

@mathmath3333 is such a wiz

OpenStudy (rational):

\(\phi(83) = 82\) and only factors of \(82\) are \(2\) and \(41\) so only possible candidates of period are : \(\{2,41,82\}\) is that how you figured out @mathmath333

OpenStudy (mathmath333):

i searched "wiz" on google

OpenStudy (mathmath333):

yes this one

OpenStudy (rational):

By similar reasoning, the period of 1/7 has to be 6 or any factors of 6 ?

OpenStudy (mathmath333):

example in \(1/73\) its period of 8 and 72=8*9 there is obvious relation with it

OpenStudy (rational):

nice nice for finding the period of \(\large \dfrac{1}{p}\), we can simply try the factors of \(\large p-1\)

myininaya (myininaya):

\[\frac{1}{43} \\ p=43 \\ p-1=42=6 \cdot 7 \] And how do you continue from here I'm still confused

myininaya (myininaya):

wait so those are the only candidates for the equation given by @mathmath333

OpenStudy (mathmath333):

\(\Large (10^{\text{try every factor of 42}} -1)\pmod {43} =0\)

OpenStudy (mathmath333):

the lowest will be the answer

myininaya (myininaya):

\[(10^x-1) (\mod p)=0 \\ (10^6-1)( \mod 43) \neq 0 \\ (10^7-1)(\mod 43) \neq 0 \\ (10^{21}-1)(\mod 43)=0\] I think I get it so the period of 1/43 is 21

OpenStudy (mathmath333):

yes

OpenStudy (rational):

thats it i guess just a side observation.. once we know that \((10^{21}-1)(\mod 43)=0\), below is also trivially true \[(10^{21x}-1)(\mod 43)=0\]

myininaya (myininaya):

That will get scarier for numbers with bigger denominators

OpenStudy (mathmath333):

not if its multiple choice question

OpenStudy (rational):

completely agree! but thats the best known method we have so far it seems..

myininaya (myininaya):

Though I guess that is the point in why they choose big numbers in encryption it runs more difficulty of us using mathematics to get the right key

OpenStudy (rational):

yeah solving \(10^x \equiv a\pmod {p}\) is a discrete logarithm problem - the most hardest problem for computers... these are used in cryptography .. solve below problem for $100,000,000 :P http://unsolvedproblems.org/index_files/RSA.htm

myininaya (myininaya):

That number is crazy and I dare @mathmath333 to take on the challenge and give me half his award :p

OpenStudy (mathmath333):

i would rather play chess tournaments and gather more

OpenStudy (mathmath333):

jk

OpenStudy (rational):

Woah thats not fair, i deserve half and you both can share 25 lol btw it is given, that number has only two factors

myininaya (myininaya):

Nothing is fair when it comes to women

myininaya (myininaya):

I do have one real question on this problem

myininaya (myininaya):

earlier to find x I was just pluggin in numbers and seeing which gave me remainder 0

myininaya (myininaya):

I guess that is the best way ?

OpenStudy (rational):

there is a small trick..

myininaya (myininaya):

or the best way we know of

myininaya (myininaya):

do tell :0

OpenStudy (mathmath333):

primitive roots ?

OpenStudy (rational):

from the fermat we know that \[10^{82}\equiv 1 \pmod{83}\]

OpenStudy (rational):

so \(10^{41}\) has to leave a remainder of \(\pm 1\) because : \[10^{82}\equiv 1 \pmod{83}\] \[10^{82}- 1 \equiv 0 \pmod{83}\] \[(10^{41}- 1)(10^{41}+1) \equiv 0 \pmod{83}\]

OpenStudy (rational):

if \(10^{41}\) leaves a remainder of \(-1\), then we're sure that the period is \(82\)

myininaya (myininaya):

that is kinda weird the period for 1/83 is (83-1)/2 and the period 1/43 is (43-1)/2

OpenStudy (rational):

the period of 1/97 is 96

OpenStudy (mathmath333):

can i substitute any primitive root of 83 for 83

myininaya (myininaya):

But it must be true for some integer n that 1/n has period (n-1)/2 and if so I wonder if we can give the set of numbers n such that is true:

OpenStudy (mathmath333):

thats too long 96

myininaya (myininaya):

1/13 has period (13-1)/2

OpenStudy (rational):

thats an interesting question primitive roots are also a good idea

OpenStudy (rational):

is \(n\) a prime in \(1/n\) ?

myininaya (myininaya):

It seems to be prime so far I'm just trying to find a set of numbers such that 1/n has period (n-1)/2

myininaya (myininaya):

but not all primes

myininaya (myininaya):

1/23 doesn't have period (23-1)/2

OpenStudy (rational):

Ohhk.. im not even sure if we can say the period function is one-to-one or one-to-many hmm

OpenStudy (mathmath333):

the (n-1)/2 will not work every time u have to check it

OpenStudy (rational):

say f(p) gives the period of 1/p in the decimal expansion where p is a prime it looks like f(p) is one-to-one...idk not so sure..

OpenStudy (rational):

can we find two primes with same period ?

myininaya (myininaya):

1/2 and 1/3 have period 1 right?

myininaya (myininaya):

or can you say 1/2 has period 1

myininaya (myininaya):

it doesn't repeat

myininaya (myininaya):

or i mean it terminates

OpenStudy (mathmath333):

i think the primitive root will not work http://www.wolframalpha.com/input/?i=10%5E%2896%29+mod+5%3D and 5 is primitive root of 97

OpenStudy (mathmath333):

2 is even it wont repeat

myininaya (myininaya):

2 is also prime

OpenStudy (rational):

Ohkk.. then there can be more than one prime having the same period

myininaya (myininaya):

so we should be looking for odd primes

OpenStudy (mathmath333):

it is the cursed prime

OpenStudy (rational):

we exclude 2 and 5 as they are factors of 10

OpenStudy (rational):

lol they are cursed primes only in base 10.. any combination of 2 and 5 in the denominator gives a terminating decimal expansion : 1/(2^a5^b) always terminates

myininaya (myininaya):

I found primes 13 and 7 and they both seem to have period 6 so it doesn't look lit it is one-to-one

myininaya (myininaya):

look like*